How many molecules are there in 4.6g of ethanol


1
Expert's answer
2021-09-10T08:00:40-0400

Molar mass of ethanol = 46.07

= 4.6/46.07 = 0.09985 mol

6.023 × 1023 × 0.09985 = 6.014 × 1022 molecules